OpenStudy (anonymous):

how do i find an equation of the line parrell to 2x + y=7 and through (5,-2)

OpenStudy (experimentx):

your equation is 3x+y = k , where k is a constant to be found since it passes through 5,-2, put the value there and find the value of k and place the value of k and you have your equation

OpenStudy (campbell_st):

the given line in poitn slope form is y = -2x + 7 the slope of the line is m = -2 then using point slope formula where, m = -2 and (5, -2) y + 2 = -2x +10 2x + y = 8

OpenStudy (experimentx):

looks like k is 8

OpenStudy (anonymous):

where did the 2 and 10 come from

OpenStudy (anonymous):

Since it's parallel, the new slope would be the same as the new slope, which is -2. Since you have your new slope and the points (5,-2) You plug into this linear equation Y-Y1=M(X-X1) M=slope=-2 Y1=-2 X1=5 And Y & X just stays like that. If you plug in the numbers, you should get. y = -2x+8